OpenStudy (anonymous):

Evaluate x/y for x =3/10 and y = 4/5

OpenStudy (luigi0210):

When dividing fractions, you have to flip them, so: \[\Huge \frac{\frac{3}{10}}{\frac{4}{5}}=\frac{3}{10}*\frac{5}{4} \]

OpenStudy (luigi0210):

Simplify by crossing out the five and ten: \[\LARGE \frac{3}{2}*\frac{1}{4}\] Multiply that

OpenStudy (isaiah.feynman):